Which pair of numbers has an even product and an odd sum?
5, 6, 9, 11

Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

5,6
5*6=30
5+6=11

Answer by checkley77(12844)

5+6=11 an odd sum.
5*6=30 an even product.
6+11=17 odd
6*11=66 even